40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Sep 19 17:02:57 2022, Paco, Summary, General, Power Outage 220916 -- restored all 
    Reply  Tue Sep 20 07:03:04 2022, Paco, Summary, General, Power Outage 220916 -- restored all 
Message ID: 17143     Entry time: Mon Sep 19 17:02:57 2022     Reply to this: 17145
Author: Paco 
Type: Summary 
Category: General 
Subject: Power Outage 220916 -- restored all 

Restore lab

[Paco, Tega, JC, Yehonathan]

We followed the instructions here. There were no major issues, apart from the fb1 ntp server sync taking long time after rebooting once.


ETMY damping

[Yehonathan, Paco]

We noticed that ETMY had to much RMS motion when the OpLevs were off. We played with it a bit and noticed two things: Cheby4 filter was on for SUS_POS and the limiter on ULCOIL was on at 0 limit. We turned both off.

We did some damping test and observed that the PIT and YAW motion were overdamped. We tune the gain of the filters in the following way:

SUSSIDE_GAIN 1250->50

SUSPOS_GAIN 200->150

SUSYAW_GAIN 60->30

These action seem to make things better.

ELOG V3.1.3-